# Energy-Aware Lane Planning for Connected Electric Vehicles in Urban Traffic: Design and Vehicle-in-the-Loop Validation

### v1

@@ I Introduction
Connected and Automated Vehicles (CAVs) provide benefits in road safety, traffic efficiency, and energy efficiency [1]. Using vehicle-to-infrastructure (V2I) and vehicle-to-vehicle (V2V) communications, CAVs can coordinate with traffic signals and neighboring vehicles to optimize their motion in ways human drivers are incapable of [2]. Prior studies have shown that by optimizing longitudinal behavior using Signal Phase and Timing (SPaT) data from connected traffic lights, a single CAV can adjust its cruising speed to avoid unnecessary stops, yielding substantial energy savings (11.35 % to 16.4%) [3, 4]. However, most existing eco-driving research [5, 6, 7] assumes that the vehicle stays in its current lane, adjusting only its longitudinal behavior. The gap is that in real-world multi-lane traffic scenarios, not only longitudinal but also lateral decisions can impact energy efficiency. For example, a CAV might save energy by changing to a lane with a “green wave” of traffic signals or to avoid following a slow vehicle [6]. This paper addresses that gap by introducing lane-changing decisions into energy-efficient motion planning for CAVs. We focus on how connectivity (V2I/V2V) can help an energy-aware lane selection strategy, extending conventional eco-driving to a multi-lane context.
In this work, we propose a novel motion planning framework that integrates lane-change decisions with energy-efficient speed optimization, enabled by V2I/V2V communication. Our approach uses a hierarchical control structure: a high-level decision module selects the most energy-efficient lane (leveraging SPaT data and V2V traffic information), and a low-level trajectory planner computes a safe and smooth trajectory to that lane. We validate the proposed system in a vehicle-in-the-loop (VIL) testing environment, embedding a real CAV into a virtual traffic simulation. This setup provides a real evaluation of energy savings and driving performance with physical vehicle dynamics and real-world like surrounding traffic, bridging the gap between pure simulation and real-world public road deployment. As with all literature on this topic, the impact of deploying this technology at scale on global energy improvement is not studied in this work.
@@ II-A Eco-Driving with V2I Connectivity
A large body of research has explored energy-efficient driving using V2I communication, particularly along signalized corridors. Optimization-based eco-driving controllers have been proposed to leverage SPaT data over corridors with multiple traffic lights. For example, in [8], the authors apply Pontryagin’s Minimum Principle to develop an eco-driving controller, demonstrating fuel savings of up to 36% compared to a human-modeled driver through VIL experiments. In [9], the authors design a real-time eco-driving controller formulated as a receding-horizon optimal control problem and demonstrate its effectiveness through real-world road testing, achieving a 31% energy reduction at the cost of increased travel time. These works demonstrate that CAVs can significantly improve efficiency through longitudinal speed planning, when fed with infrastructure information. However, these studies assume the vehicle stays in its lane and controls speed only. The controllers react to traffic slowdowns or red lights by decelerating or coasting, but do not consider proactive lane changes to avoid those slowdowns.
@@ II-B Lane Changes in Eco-Driving
Researchers, only recently, have begun to consider lateral maneuvers as part of the eco-driving problem. It is well understood that getting stuck behind a slow or stopping vehicle can negate the benefits of an optimal speed profile. In [10], the authors explicitly note that many eco-driving studies focus on longitudinal control and overlook lane changes, which may impair fuel efficiency when a slower vehicle is ahead. To address this, the authors propose a high-level graph search that selects the optimal lane sequence based on predicted traffic in each lane. However, the approach does not account for traffic lights or their influence on their strategy. Similarly, in [11], the authors introduce an overtaking strategy: they first compute an energy-optimal speed trajectory for a given route, then allow the CAV to make a lane change if a slower vehicle ahead would disturb that trajectory. These studies collectively demonstrate that lane-changing can complement longitudinal eco-driving. However, these lane-changing eco-driving strategies so far have been evaluated only in simulation. Little work has been reported on the implementation of such strategies in real vehicles or even high-fidelity VIL setups, except our previous work [12].
@@ III Contributions
This paper makes the following contributions to energy-efficient motion planning for connected, automated urban driving:
  • Formulates the lane selection problem in urban environments as a behavior planning task that jointly considers both lateral and longitudinal actions under SPaT and surrounding traffic aware constraints.
  • Proposes a hybrid lane selection algorithm that solves short-horizon optimization problems and approximates the long-horizon energy costs using a graph search-based terminal cost-to-go.
  • Incorporates a data-driven energy regression model calibrated to the Hyundai IONIQ5, allowing quantitative evaluation of energy consumption for different lane and speed choices.
  • Demonstrates, through vehicle-in-the-loop experiments, that the proposed method reduces motion energy consumption by 24% compared to human driving, and achieves approximately 7% reduction compared to a no-lane-change eco-driving baseline in the scenario studied in this paper.
@@ IV-A Lane Selection Problem
This paper focuses on a two-lane, one-way road, as illustrated in Fig. 2, although the proposed methods can be extended to roads with more than two lanes. We assume that the CAV can perfectly observe surrounding vehicles’ states within a certain radius and, by leveraging vehicle-to-infrastructure (V2I) connectivity, can receive SPaT information for the next $N_{TL}$ traffic lights along its route. The available information to the lane selector includes the controlled vehicle’s state, denoted as $x^{EV}(t)=[s(t),v(t)]^{\top}$, where $s(t)$ and $v(t)$ represent the longitudinal displacement along the centerline of the given route and the longitudinal velocity, respectively. Additionally, the vehicle’s lane information is encoded as a categorical variable $lane(t)\in\{0,1\}$. The surrounding vehicles (SV) within the detection range are collected into
$o(t)=[x^{i}(t),lane^{i}(t)]_{i=1}^{N_{SV}}.$ (1)
where $N_{SV}$ is the number of detected SVs. Lastly, the SPaT information of the next $N_{TL}$ upcoming traffic lights is collected into
@@
where $s_{tl}^{i}\in\mathbb{R}$, $p_{tl}^{i}(t)\in\{r,g,y\}$, $t_{tl}^{i}(t)\in\mathbb{R}$ denote the longitudinal position of the $i$-th traffic light along the centerline of the route, the current signal phase (red, green, or yellow), and the remaining time in the current phase, respectively. The duration of the full cycle of traffic signals—green ($t^{i}_{g}$), yellow ($t^{i}_{y}$), and red ($t^{i}_{r}$)—are assumed to be estimated from historic SPaT data. For brevity, let $\mathcal{C}_{\text{pass}}$ denote the lane selector’s decision regarding whether the controlled vehicle will pass the upcoming traffic light and which lane it should proceed in. In the case of two lanes,
[…]
Now, the succinct lane selection problem is to design a mapping function
$\mathcal{C}_{\text{pass},t}=f(x^{\text{EV}}(t),o(t),SPaT(t)),$ (4)
@@ IV-B1 Data-driven Energy Consumption Model [7]
We approximate the energy consumption of the battery electric vehicle (BEV) given a longitudinal trajectory. Specifically, we assume that the energy consumption is a function of the longitudinal displacement $\mathbf{s}$ along the centerline and the longitudinal acceleration $\mathbf{a}$ of the vehicle. For more details, refer to Sec. II.C. in [7]. By recording the vehicle’s state-of-charge (SOC) and longitudinal motion, we fit a quadratic regression model to estimate the energy consumption rate $\Delta E$, denoted as $\mathcal{E}(v,a)$, where $v$ and $a$ are the vehicle’s longitudinal velocity and acceleration. The quadratic regression model is parameterized as
$\mathcal{E}(v,a)=z^{\top}Pz+q^{\top}z+r$ (5)
@@
$\mathcal{C}^{\text{eco}}_{\text{pass},t}=f(x^{\text{EV}}(t),o(t),SPaT(t),\mathcal{E}).$ (6)
@@ V-A1 Vehicle Model
where $u(t)=a(t)$ which is the longitudinal acceleration of the vehicle. We discretize the dynamics (7) using the sampling time $\Delta t$ of 0.5 seconds. The discretized dynamics are denoted as
@@ V-A2 Graph Search Terminal Cost Approximation
[…]
@@
where $v(V)$ represents the passing velocity value of the node and $\ell$ denotes the $\ell$-th traffic light that the node $V_{j}$ represents. $tl^{\ell,p}$ at the time the ego vehicle reaches the traffic light is computed based on the travel time between nodes assuming the vehicle travels at a constant velocity between traffic lights. $\mathbb{1}(\cdot)$ is the Heaviside step function. The connections from nodes in $\mathcal{V}_{N_{TL}}$ to the terminal node $V_{-1}$ are added to penalize the velocity deviation from the velocity at the point of passing the first traffic light. Alternatively, the data-driven energy model in (5) can be used to approximate the energy consumption resulting from changes in node velocities, assuming constant acceleration. However, for simplicity, this work adopts the kinetic energy model to compute edge costs. Note that the graph (9) satisfies additive property meaning
$c(V_{a},V_{c})=c(V_{a},V_{b})+c(V_{b},V_{c})$ (11)
for any node connections $V_{a}\rightarrow V_{b}$ and $V_{b}\rightarrow V_{c}$. Additivity allows us to exploit dynamic programming and optimal substructure, formulating the problem as a path-finding problem and applying Dijkstra’s algorithm [14] to compute $\min c(V_{1},V_{-1})$ given $v(V_{1})$. We define the minimum sum of the edge cost from the starting node to the terminal node as
@@
and use it to approximate the energy cost from the 1st traffic light to the $N_{TL}$-th traffic light assuming the ego vehicle travels in the fixed lane and at a constant velocity.
@@ V-A3 Constraints
$\mathcal{X}$ :=$:=\{x\in\mathbb{R}^{2}|\ 2\leq v\leq 11,\text{where}\;x=[s,v]^{\top}\},$ (13)
$\mathcal{U}$ :=$:=\{u\in\mathbb{R}|-4\leq u\leq 2\}.$ (14)
@@
[…]
$\text{Cycle}:=[t^{\text{next}}_{g},t^{\text{next}}_{y},t^{\text{next}}_{r}],$ (18)
where $t^{\text{next}}_{g}$, $t^{\text{next}}_{y}$, and $t^{\text{next}}_{r}$ denote the durations of the green, yellow, and red phases respectively—which can be obtained from the traffic light’s SPaT data—the $CT$ function is defined as follows:
$CT(\phi,\mathcal{C}_{\text{pass}}):=t_{tl}^{\text{next}}(t)+\sum_{i=1}^{\delta(\phi,\mathcal{C}_{\text{pass}})}\text{Cycle}[(\phi+i)\bmod 3],$ (19)
@@ V-A4 Optimal Control Problem
We solve the following constrained finite-time optimal control problem for every $\mathcal{C}_{\text{pass}}$ candidates in the set (3):
$J$ (${}^{\star}_{\text{ftocp}}(x(t),u(t))=$
@@
$\quad x_{k|t}\in\mathcal{X},$ (21c)
$\quad u_{k|t}\in\mathcal{U},$ (21d)
$\quad(x_{k|t},u_{k|t})\in\mathcal{F}^{LK}(s^{\text{front}}_{k|t},v^{\text{front}}_{k|t})$ (21e)
$\quad x_{t|t}=x(t),u_{t|t}=u(t),\leavevmode\nobreak\ \forall k\in\mathbb{I}_{t}^{t+N-1}$ (21f)
$\quad x_{k|t}\in\mathcal{P}(tl^{\text{next},s}),\leavevmode\nobreak\ \forall k\in\mathbb{I}_{[t+\underline{t}_{\text{pass}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}^{[t+\bar{t}_{\text{pass}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}$ (21g)
$\quad x_{k|t}\in\mathcal{S}(tl^{\text{next},s}),\leavevmode\nobreak\ \forall k\in\mathbb{I}_{[t+\underline{t}_{\text{stop}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}^{[t+\bar{t}_{\text{stop}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}$ (21h)
@@
$\mathcal{C}^{\text{eco}}_{\text{pass}}=\operatorname*{argmin}_{\mathcal{C}\in\mathcal{LS}}\;J^{\star}(\mathcal{C})$ (23)
The lane selector also provides the trajectory planner with the reference velocity profile $v_{\text{ref}}$, which represents the energy-efficient speed trajectory corresponding to the selected behavior decision $\mathcal{C}^{\text{eco}}_{\text{pass}}$. The lower-level trajectory planner then computes a motion trajectory that closely tracks this reference velocity while respecting traffic laws and ensuring safety with respect to surrounding NPC vehicles, using a finer time discretization to guarantee dynamic feasibility and reactivity [12].

### v2

@@ I Introduction
Connected and Automated Vehicles (CAVs) can improve safety, traffic flow, and energy efficiency by coordinating with traffic signals and nearby vehicles through vehicle-to-infrastructure (V2I) and vehicle-to-vehicle (V2V) communication [1]. Eco-driving strategies leveraging Signal Phase and Timing (SPaT) data from connected traffic lights have demonstrated 11–16% energy savings [2, 3], but these methods typically assume fixed-lane driving and optimize only longitudinal motion [4, 5, 6]. In practice, lane changes also influence efficiency by avoiding slower vehicles or aligning with favorable traffic signals [5]. Yet most eco-driving studies overlook this factor, limiting their applicability to realistic multi-lane scenarios. We extend eco-driving to include lane-change decisions for energy savings for CAVs. We focus on how connectivity (V2I/V2V) can help an energy-aware lane selection strategy, extending conventional eco-driving to a multi-lane context.
In this work, we propose a novel motion planning framework that integrates lane-change decisions with energy-efficient speed optimization, enabled by V2I/V2V communication. Our approach uses a hierarchical control structure: a high-level decision module selects the most energy-efficient lane (leveraging SPaT data and V2V traffic information), and a low-level trajectory planner computes a safe and smooth trajectory to that lane. We validate the proposed system in a vehicle-in-the-loop (VIL) testing environment, embedding a real CAV into a virtual traffic simulation. This setup provides a real evaluation of energy savings and driving performance with physical vehicle dynamics and realistic surrounding traffic, bridging the gap between pure simulation and real-world public road deployment. As with all literature on this topic, the impact of deploying this technology at scale on global energy improvement is not studied in this work.
@@ II-A Eco-Driving with V2I Connectivity
A large body of research has explored energy-efficient driving using V2I communication, particularly along signalized corridors. Optimization-based eco-driving controllers have been proposed to leverage SPaT data over corridors with multiple traffic lights. For example, in [7], the authors apply Pontryagin’s Minimum Principle to develop an eco-driving controller, demonstrating fuel savings of up to 36% compared to a human-modeled driver through VIL experiments. In [8], the authors design a real-time eco-driving controller formulated as a receding-horizon optimal control problem and demonstrate its effectiveness through real-world road testing, achieving a 31% energy reduction at the cost of increased travel time. These works demonstrate that CAVs can significantly improve efficiency through longitudinal speed planning, when leveraging infrastructure information. However, these studies assume a fixed lane. The controllers react to surrounding traffics and red lights, but do not consider proactive lane changes to avoid deceleration.
@@ II-B Lane Changes in Eco-Driving
Only recently have researchers begun to incorporate lateral maneuvers into eco-driving. In [9], the authors note that most studies emphasize longitudinal control and neglect lane changes, which can reduce efficiency when following slower vehicles. To address this, they propose a graph-search strategy that selects lane sequences based on predicted traffic, though the effect of traffic lights is not considered. Similarly, [10] designs an overtaking strategy in which an energy-optimal speed trajectory is first computed, and a lane change is triggered if a slower vehicle disrupts that plan. These studies indicate that lane-changing can enhance longitudinal eco-driving, yet their validation remains limited to simulation. Few works have demonstrated lane-changing eco-driving in real vehicles or high-fidelity VIL platforms, aside from our previous work [11].
@@ III Contributions
This work’s contributions to energy-efficient motion planning for connected, automated urban driving are as follows
  • Formulation of the lane selection problem in urban environments as a joint behavior planning task over lateral and longitudinal actions, using SPaT and surrounding traffic information.
  • A hybrid lane selection algorithm that combines short-horizon optimal control with a graph-based approximation of long-term energy costs, supported by a data-driven energy model calibrated on a Hyundai IONIQ5.
  • Vehicle-in-the-loop experiments demonstrating up to 24% motion energy reduction compared to human driving and 7% savings over a lane-keeping eco-driving baseline.
@@ IV-A Lane Selection Problem
This paper focuses on a two-lane, one-way road, as illustrated in Fig. 2, although the proposed methods can be extended to roads with more than two lanes. We assume that the CAV can perfectly observe surrounding vehicles’ states within a certain radius and, by leveraging V2I connectivity, can receive SPaT information for the next $N_{TL}$ traffic lights along its route. The information available to the lane selector includes the controlled vehicle’s state, denoted as $x^{EV}(t)=[s(t),v(t)]^{\top}$, where $s(t)$ and $v(t)$ are the longitudinal displacement along the centerline of the given route and the longitudinal velocity, respectively. Additionally, the vehicle’s lane information is encoded as a categorical variable $lane(t)\in\{0,1\}$. The surrounding vehicles (SV) within the detection range are collected into
$o(t)=[x^{i}(t),lane^{i}(t)]_{i=1}^{N_{SV}},$ (1)
where $N_{SV}$ is the number of detected SVs. Lastly, the SPaT of the next $N_{TL}$ upcoming traffic lights is collected into
@@
where $s_{tl}^{i}\in\mathbb{R}$, $p_{tl}^{i}(t)\in\{r,g,y\}$, $t_{tl}^{i}(t)\in\mathbb{R}$ denote the longitudinal position of the $i$-th traffic light along the centerline of the route, the current signal phase (red, green, or yellow), and the remaining time in the current phase, respectively. The duration of the full cycle of traffic signals—green ($t^{i}_{g}$), yellow ($t^{i}_{y}$), and red ($t^{i}_{r}$)—are assumed to be estimated from historic SPaT data. Let $\mathcal{C}_{\text{pass}}$ denote the lane selector’s decision regarding whether the controlled vehicle will pass the upcoming traffic light and which lane it should proceed in. In the case of two lanes,
$\mathcal{C}_{\text{\scriptsize pass}}\in\mathcal{LS}:=\{\text{\scriptsize PASS0},\text{\scriptsize PASS1},\text{\scriptsize NONPASS0},\text{\scriptsize NONPASS1}\},$ (3)
where PASS0 and PASS1 denote the lane selector’s decision to pass the upcoming traffic light in lane 0 and lane 1, respectively. NONPASS0 and NONPASS1 denote the decision to not pass the upcoming traffic light in lane 0 and lane 1, respectively. The subset of passing decisions is denoted by $\mathcal{LS}_{\text{PASS}}:=\{\text{PASS0},\text{PASS1}\}$, and the subset of nonpassing decisions is denoted by $\mathcal{LS}_{\text{NONPASS}}:=\{\text{NONPASS0},\text{NONPASS1}\}$.
Now, the lane selection problem is to design a mapping function
$\small\mathcal{C}_{\text{pass},t}=f(x^{\text{EV}}(t),o(t),SPaT(t)),$ (4)
@@ IV-B1 Data-driven Energy Consumption Model [6]
We approximate the energy consumption of the battery electric vehicle (BEV) given a longitudinal trajectory. Specifically, we assume that the energy consumption is a function of the longitudinal displacement along the centerline and the longitudinal acceleration of the vehicle. For more details, refer to Sec. II.C. in [6]. By recording the vehicle’s state-of-charge (SOC) and longitudinal motion, we fit a quadratic regression model to estimate the energy consumption rate $\Delta E$, denoted as $\mathcal{E}(v,a)$, where $v$ and $a$ are the vehicle’s longitudinal velocity and acceleration. The quadratic regression model is parameterized as
$\small\mathcal{E}(v,a)=z^{\top}Pz+q^{\top}z+r$ (5)
@@
$\small{\mathcal{C}^{\text{eco}}_{\text{pass},t}=f(x^{\text{EV}}(t),o(t),SPaT(t),\mathcal{E}).}$ (6)
@@ V-A1 Vehicle Model
where $u(t)=a(t)$ is the longitudinal acceleration of the vehicle. We discretize the dynamics (7) using the sampling time $\Delta t$ of 0.5 seconds. The discretized dynamics are denoted as
@@ V-A2 Graph Search Terminal Cost Approximation
where $V=\{V_{1},\mathcal{V}_{2},\dots,\mathcal{V}_{N_{TL}},V_{-1}\}$ and $\mathcal{V}_{i}$, $\forall i\in\mathbb{I}_{1}^{N_{TL}}$, represents the subset of nodes corresponding to each traffic light as depicted in Fig. 4. Each node in $\mathcal{V}_{i}$ represents the velocity at the point of passing the $i$-th traffic light. We grid the passing velocities from $v^{G}_{min}=5\,\text{m/s}$ to $v^{G}_{max}=10\,\text{m/s}$ with a grid size of $1\,\text{m/s}$. The starting and terminal nodes $V_{1}$ and $V_{-1}$ represent the first traffic light ahead and the ego vehicle’s position after passing the $N_{TL}$-th traffic light, respectively. The solution of the optimal control problem determines the passing velocity at the 1st traffic light. For simplicity, the passing velocity at the terminal node is set equal to that of the starting node.
@@
where $v(V)$ represents the passing velocity value of the node and $\ell$ denotes the $\ell$-th traffic light that the node $V_{j}$ represents. $tl^{\ell,p}$ at the time the ego vehicle reaches the traffic light is computed based on the travel time between nodes assuming the vehicle travels at a constant velocity between traffic lights. $\mathbb{1}(\cdot)$ is the Heaviside step function. The connections from nodes in $\mathcal{V}_{N_{TL}}$ to the terminal node $V_{-1}$ are added to penalize the velocity deviation from the velocity at the point of passing the first traffic light. Alternatively, the data-driven energy model in (5) can be used to approximate the energy consumption resulting from changes in node velocities, assuming constant acceleration. However, for simplicity, this work adopts the kinetic energy model to compute edge costs. Note that the graph (9) satisfies additive property meaning $c(V_{a},V_{c})=c(V_{a},V_{b})+c(V_{b},V_{c})$ for any node connections $V_{a}\rightarrow V_{b}$ and $V_{b}\rightarrow V_{c}$. Additivity allows us to exploit dynamic programming and optimal substructure, formulating the problem as a path-finding problem and applying Dijkstra’s algorithm [13] to compute $\min c(V_{1},V_{-1})$ given $v(V_{1})$. We define the minimum sum of the edge cost from the starting node to the terminal node as
@@
and use it to approximate the energy cost from the 1st traffic light to the $N_{TL}$-th traffic light, assuming the ego vehicle travels in the fixed lane and at a constant velocity.
@@ V-A3 Constraints
$\mathcal{X}$ $:=\{x\in\mathbb{R}^{2}|\ 2\leq v\leq 11,\text{where}\;x=[s,v]^{\top}\},$ (12)
$\mathcal{U}$ $:=\{u\in\mathbb{R}|-4\leq u\leq 2\}.$ (13)
@@
where $\bar{t}_{\text{pass}}$ and $\underline{t}_{\text{pass}}$ denote the upper and lower bounds of the time interval during which the pass constraint is active; analogous definitions apply to the stop constraint. The parameter $t_{g}^{\text{margin}}$ is a user-defined safety margin that ensures the vehicle passes through the traffic light within the green phase. In this work, we set $t_{g}^{\text{margin}}=2$ s. Additionally, the function $CT$ denotes the cycle time adjustment, which accounts for the current signal phase of the upcoming traffic light and the decision to either pass or stop. Let $\phi\in\{0,1,2\}$ be the categorical encoding of the next traffic light’s signal phase, corresponding to green, yellow, and red respectively: $0\mapsto g,\ 1\mapsto y,\ 2\mapsto r]$. The next traffic light’s cycle durations is $\text{Cycle}:=[t^{\text{next}}_{g},t^{\text{next}}_{y},t^{\text{next}}_{r}],$ where $t^{\text{next}}_{g}$, $t^{\text{next}}_{y}$, and $t^{\text{next}}_{r}$ denote the durations of the green, yellow, and red phases respectively—which can be obtained from the traffic light’s SPaT data. The $CT$ function is defined as follows:
$CT(\phi,\mathcal{C}_{\text{pass}}):=t_{tl}^{\text{next}}(t)+\hskip-6.00006pt\sum_{i=1}^{\delta(\phi,\mathcal{C}_{\text{pass}})}\hskip-6.00006pt\text{Cycle}[(\phi+i)\bmod 3],$ (17)
@@ V-A4 Optimal Control Problem
We solve the following constrained finite-time optimal control problem for every $\mathcal{C}_{\text{pass}}$ candidates in the set (3) and collect the optimal cost $J^{\star}_{\text{ftocp}}(x(t),u(t))$:
@@
$\quad x_{k|t}\in\mathcal{X},u_{k|t}\in\mathcal{U},$ (19c)
$\quad(x_{k|t},u_{k|t})\in\mathcal{F}^{LK}(s^{\text{front}}_{k|t},v^{\text{front}}_{k|t})$ (19d)
$\quad x_{t|t}=x(t),u_{t|t}=u(t),\leavevmode\nobreak\ \forall k\in\mathbb{I}_{t}^{t+N-1}$ (19e)
$\quad x_{k|t}\in\mathcal{P}(tl^{\text{next},s}),\leavevmode\nobreak\ \forall k\in\mathbb{I}_{[t+\underline{t}_{\text{pass}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}^{[t+\bar{t}_{\text{pass}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}$ (19f)
$\quad x_{k|t}\in\mathcal{S}(tl^{\text{next},s}),\leavevmode\nobreak\ \forall k\in\mathbb{I}_{[t+\underline{t}_{\text{stop}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}^{[t+\bar{t}_{\text{stop}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}$ (19g)
@@
$\vskip-2.84544pt\mathcal{C}^{\text{eco}}_{\text{pass}}=\operatorname*{argmin}_{\mathcal{C}\in\mathcal{LS}}\;J^{\star}(\mathcal{C})$ (21)
The lane selector also provides the reference velocity profile $v_{\text{ref}}$, which represents the energy-efficient speed trajectory corresponding to the selected behavior decision $\mathcal{C}^{\text{eco}}_{\text{pass}}$. The lower-level trajectory planner then computes a motion trajectory that closely tracks this reference velocity while respecting traffic laws and ensuring safety with respect to surrounding vehicles, using a finer time discretization to guarantee dynamic feasibility and reactivity [11].